### v4

# As Soon as Possible but Rationally

Abstract
This paper addresses complexity problems in rational verification and synthesis for multi-player games played on weighted graphs, where the objective of each player is to minimize the cost of reaching a specific set of target vertices. In these games, one player, referred to as the system, declares his strategy upfront. The other players, composing the environment, then rationally make their moves according to their objectives. The rational behavior of these responding players is captured through two models: they opt for strategies that either represent a Nash equilibrium or lead to a play with a Pareto-optimal cost tuple.
keywords:
Games played on graphs, rational verification, rational synthesis, Nash equilibrium, Pareto-optimality, quantitative reachability objectives
1 Introduction
Nowadays, formal methods play a crucial role in ensuring the reliability of critical computer systems. Still, the application of formal methods on a large scale remains elusive in certain areas, notably in multi-agent systems. Those systems pose a significant challenge for formal verification and automatic synthesis because of their heterogeneous nature, encompassing everything from conventional reactive code segments to fully autonomous robots and even human operators. Crafting formal models that accurately represent the varied components within these systems is often a too complex task.
Although constructing detailed operational models for humans or sophisticated autonomous robots might be problematic, it is often more feasible to identify the overarching goals that those agents pursue. Incorporating these goals is crucial in the design and validation process of systems that interact with such entities. Typically, a system is not expected to function flawlessly under all conditions but rather in scenarios where the agents it interacts with act in alignment with their objectives, i.e., they behave rationally. Rational synthesis focuses on creating a system that meets its specifications against any behavior of environmental agents that is guided by their goals (and not against any of their behaviors). Rational verification studies the problem of ensuring that a system enforces certain correctness properties, not in all conceivable scenarios, but specifically in scenarios where environmental agents behave in accordance with their objectives.
Rationality can be modeled in various ways. In this paper, we focus on two general approaches. The first approach comes from game theory where rationality is modeled by the concept of equilibrium, such as Nash equilibria (NE) [Nash50] or subgame perfect equilibria (SPE), a refinement of NEs [osbornebook]. The second approach treats the environment as a single agent but with multiple, sometimes conflicting, goals, aiming for behaviors that achieve a Pareto-optimal balance among these objectives. The concept of Pareto-optimality (PO) and its application in multi-objective analysis have been explored primarily in the field of optimization [DBLP:conf/focs/PapadimitriouY00], but also in formal methods [DBLP:conf/fossacs/AlurDMW09, DBLP:conf/cav/BrenguierR15]. These two notions of rationality are different in nature: in the first setting, each component of the environment playing an equilibrium is considered to be an independent selfish individual, excluding cooperation scenarios, while in the second setting, several components of the environment can collaborate and agree on trade-offs. The challenge lies in adapting these concepts to reactive systems characterized by ongoing, non-terminating interactions with their environment. This necessitates the transition from two-player zero-sum games on graphs, the classical approach used to model a fully adversarial environment (see e.g. [PnueliR89]), to the more complex setting of multi-player non zero-sum games on graphs used to model environments composed of various rational agents.
Rational synthesis and rational verification have attracted large attention recently, see e.g. [DBLP:conf/mfcs/BriceRB23, Pareto-Rational-Verification, ConduracheFGR16, FismanKL10-crsp-kupferman, DBLP:journals/ai/GutierrezNPW20, DBLP:journals/amai/GutierrezNPW23, KupfermanPV14-ncrsp, tacas2022]. But the results obtained so far, with a few exceptions that we detail below, are limited to the qualitative setting formalized as Boolean outcomes associated with $\omega$-regular objectives. Those objectives are either specified using linear temporal specifications or automata over infinite words (like parity automata). The complexity of those problems is now well understood (with only a few complexity gaps remaining, see e.g. [ConduracheFGR16, tacas2022]). The methods to solve those problems and get completeness results for worst-case complexity are either based on automata theory (using mainly automata over infinite trees) or by reduction to zero-sum games. Quantitative objectives are less studied and revealed to be much more challenging. For instance, it is only very recently that the rational verification problem was studied, for SPEs in non zero-sum games with mean-payoff, energy, and discounted-sum objectives in [DBLP:conf/mfcs/BriceRB23], for an LTL specification in multi-agent systems that behave according to an NE with mean-payoff objectives in [DBLP:journals/amai/GutierrezNPW23] or with quantitative probabilistic LTL objectives in [DBLP:conf/atal/Hyland0KW24]. In [AlmagorKP18], the rational synthesis problem was studied for the quantitative extension $\mathcal{F}$] of LTL where the Boolean operators are replaced with arbitrary functions mapping binary tuples into the interval $[0,1]$.
In this paper, we consider quantitative reachability objectives. Our choice for studying these objectives was guided by their fundamental nature and also by their relative simplicity. Nevertheless, as we will see, they are challenging for both rational synthesis and rational verification. Indeed, to obtain worst-case optimal algorithms and establish completeness results, we had to resort to the use of innovative theoretical tools, more advanced than those necessary for the qualitative framework. In our endeavor, we have established the exact complexity of most studied decision problems in rational synthesis and rational verification.
[b]
Non-coop. verif. Universal non-coop. verif. Coop. synthesis Non-coop. synthesis
PO, weights $\Pi_{2}^{\mathsf{P}}$-complete $\mathsf{PSPACE}$-complete $\mathsf{PSPACE}$-complete $\mathsf{NEXPTIME}$-complete [gaspard-pareto-quanti-reach]
PO, qualitative $\Pi_{2}^{\mathsf{P}}$-complete $\mathsf{PSPACE}$-complete $\mathsf{PSPACE}$-complete $\mathsf{NEXPTIME}$-complete [Stackelberg-Pareto-Synthesis]
NE, weights $\mathsf{coNP}$-complete $\mathsf{coNP}$-complete $\mathsf{NP}$-complete Unknown, $\mathsf{EXPTIME}$-hard1
NE, qualitative $\mathsf{coNP}$-complete [grandmont23] $\mathsf{coNP}$-complete [grandmont23] $\mathsf{NP}$-complete [ConduracheFGR16] $\mathsf{PSPACE}$-complete [ConduracheFGR16]
  • 1
    For the important special case of one-player environments, we provide an algorithm that runs in $\mathsf{EXPTIME}$ and we can prove $\mathsf{PSPACE}$-hardness. The $\mathsf{EXPTIME}$-hardness of the general case already holds for two-player environments.
Table 1: Summary of complexity results.
Technical Contributions.
In this work, we explore both verification and synthesis problems through the lens of rationality, defined by Pareto-optimality and Nash equilibria, for quantitative reachability objectives. For the synthesis problem, we also consider the cooperative variant where the environment cooperates with the system: we want to decide whether the system has a strategy and the environment a rational response to this strategy such that the objective of the system is enforced. Our results are presented in Table 1, noting that all results lacking explicit references are, to our knowledge, novel contributions. For completeness, the table includes (new and known) results for the qualitative scenario.
The results for PO rationality are as follows. (1) For the verification problems, we assume that the behavior of the system is formalized by a nondeterministic Mealy machine, used to represent a (usually infinite) set of its possible implementations. For each of those implementations, we verify that the quantitative reachability objective of the system is met against any rational behavior of the environment. We establish that this problem is $\mathsf{PSPACE}$-complete. To obtain the upper bound, we rely on a genuine combination of techniques based on Parikh automata and a recursive $\mathsf{PSPACE}$ algorithm (for positive Boolean combinations of bounded safety objectives, a problem of independent interest). Parikh automata are used to guess a compact representation of certificates which are paths of possibly exponential length in the size of the problem input. When the Mealy machine is deterministic, we show that the complexity goes down to $\Pi_{2}^{\mathsf{P}}$-completeness, as the previous $\mathsf{PSPACE}$ algorithm is replaced by a $\mathsf{coNP}$ oracle. (2) For the synthesis problems, we only consider the cooperative version which we prove to be $\mathsf{PSPACE}$-complete, as the non-cooperative version was established to be $\mathsf{NEXPTIME}$-complete in [gaspard-pareto-quanti-reach].
The results for NE rationality are as follows. (1) We establish that, surprisingly, the verification problems are $\mathsf{coNP}$-complete both for the general case of a nondeterministic Mealy machine and for the special case where it is deterministic. The upper bounds for those problems are again based on Parikh automata certificates but here there is no need to use a $\mathsf{coNP}$ oracle. (2) For the synthesis problems, the landscape is more challenging. For the cooperative case, we were able to establish that the problem is $\mathsf{NP}$-complete. For the non-cooperative case, we have partially solved the problem and established the following results. When the environment is composed of a single rational player, the problem is in $\mathsf{EXPTIME}$ and $\mathsf{PSPACE}$-hard. For an environment with at least two players, we show that the problem is $\mathsf{EXPTIME}$-hard but we leave its decidability open. The lower bounds are obtained using an elegant reduction from countdown games [countdown-game]. We give indications in the paper why the problem is difficult to solve and why classical automata-theoretic methods may not be sufficient (if the problem is decidable).
In this paper, we focus on nonnegative weights as we show that considering arbitrary weights leads to undecidability of the synthesis problems. We also focus on NEs instead of SPEs, even if the latter are a better concept to model rational behavior in games played on graphs. Indeed, it is well-known that SPEs pose greater challenges than NEs. So, starting with NEs offers a better initial step for the algorithmic treatment of rational verification and synthesis in quantitative scenarios, an area that remains largely unexplored.
Related Work.
The survey [DBLP:conf/fsttcs/BrihayeGMR23] presents several results about different game models and different kinds of objectives related to reachability. Quantitative objectives in two-player zero-sum games were largely studied, see e.g. [DBLP:conf/concur/BrihayeGHM15, DBLP:conf/fsttcs/ChatterjeeDHR10, ehrenfeucht1979positional], even if exact complexity results are often elusive due to the intricate nature of the problems (e.g. the exact complexity of solving mean-payoff games is still an open problem). In multi-player non zero-sum games, the (constrained) existence of equilibria is also well studied. The existence of simple NEs was established in [DBLP:conf/lfcs/BrihayePS13] for mean-payoff and discounted-sum objectives. No decision problem is considered in that paper. The constrained existence of SPEs in quantitative reachability games was proved $\mathsf{PSPACE}$-complete in [DBLP:journals/iandc/BrihayeBGR21]. We prove here that the complexity is lower when we use NEs to model rationality, as we obtain $\mathsf{NP}$-completeness for the related cooperative synthesis problem. Deciding the constrained existence of SPEs was recently solved for quantitative reachability games in [DBLP:conf/concur/BrihayeBGRB19] and for mean-payoff games in [DBLP:conf/concur/BriceRB21, DBLP:conf/icalp/BriceRB22]. The cooperative and non-cooperative rational synthesis problems were studied in [NoncoopSynth_Meanpayoff_2players] for games with mean-payoff and discounted-sum objectives when the environment is composed of a single player. The mean-payoff case was proved to be $\mathsf{NP}$-complete and the discounted-sum case was linked to the open target discounted sum problem, which explains the difficulty of solving the problem in this case.
Structure of the Paper.
The background is given in Section 2. The formal definitions of the studied problems and our main complexity results are stated in Section 3. The proofs of our results are given for PO rationality in LABEL:section:pareto, and for NE rationality in LABEL:section:Nash. We give a conclusion and future work in LABEL:sec:conclusion.
2 Background
Arenas and Plays.
A (finite) arena $\mathcal{A}{}$ is a tuple $(V,E,\mathcal{P},(V_{i})_{i\in\mathcal{P}})$ where $V$ is a finite set of vertices, $E\subseteq V\times V$ is a set of edges, $\mathcal{P}$ is a finite set of players, and $(V_{i})_{i\in\mathcal{P}}$ is a partition of $V$, where $V_{i}$ is the set of vertices owned by player $i$. We assume that $v\in V$ has at least one successor, i.e., the set $\mathsf{succ}({v})=\{v^{\prime}\in V\mid(v,v^{\prime})\in E\}$ is nonempty.
[…]
We can concatenate two nonempty histories $h_{1}$ and $h_{2}$ into a single one, denoted $h_{1}\cdot h_{2}$ or $h_{1}h_{2}$ if $(\mathsf{last}({h_{1}}),\mathsf{first}({h_{2}}))\in E$. When a history can be concatenated to itself, we call it a cycle. Furthermore, a play $\pi=\mu\nu\nu\dots=\mu(\nu)^{\omega}$ where $\mu\nu\in\mathsf{Hist}$ with $\nu$ a cycle, is called a lasso. The length of $\pi$ is then the length of $\mu\nu$ [^2] To have a well-defined length for a lasso $\pi$, we assume that $\pi=\mu(\nu)^{\omega}$ with $\mu\nu$ of minimal length. Given a play $\pi$, a cycle along $\pi$ refers to a sequence $\pi_{[m,n[}$ with $\pi_{m}=\pi_{n}$. We denote $\mathsf{Occ}({\pi})=\{v\in V\mid\exists n\in\mathbb{N},\;v=\pi_{n}\}$ the set of all vertices occurring along $\pi$, and we say that $\pi$ visits or reaches a vertex $v\in\mathsf{Occ}({\pi})$ or a set $T$ such that $T\cap\mathsf{Occ}({\pi})\neq\emptyset$. The previous notions extend to histories.
Given an arena $\mathcal{A}$, if we fix an initial vertex $v_{0}\in V$, we say that $\mathcal{A}$ is initialized and we denote by $\mathsf{Plays}(v_{0})$ (resp. $\mathsf{Hist}(v_{0})$) all its plays (resp. nonempty histories) starting with $v_{0}$. An arena is called weighted if it is augmented with a non-negative weight function $w_{i}:E\rightarrow\mathbb{N}$ for each player $i$. We denote by $W$ the greatest weight, i.e., $W=\max\{w_{i}(e)\mid e\in E,i\in\mathcal{P}\}$. We extend $w_{i}$ to any history $h=\pi_{0}\dots\pi_{n}$ such that $w_{i}(h)=\sum_{j=1}^{n}\,w_{i}((\pi_{j-1},\pi_{j}))$.
Reachability Games.
A reachability game is a tuple $\mathcal{G}{}=(\mathcal{A},(T_{i})_{i\in\mathcal{P}})$ where $\mathcal{A}$ is a weighted arena and $T_{i}\subseteq V$ is a target set for each $i\in\mathcal{P}$. We define a cost function $\mathsf{cost}_{i}:\mathsf{Plays}\rightarrow\mathbb{N}\cup\{+\infty\}$ for each player $i$, such that for all plays $\pi=\pi_{0}\pi_{1}\dots\in\mathsf{Plays}$, $\mathsf{cost}_{i}(\pi)=w_{i}(\pi_{0}\dots\pi_{n})$ with $n$ the smallest index such that $\pi_{n}\in T_{i}$, if it exists and $\mathsf{cost}_{i}(\pi)=+\infty$ otherwise.
The reachability objective of player $i$ is to minimize this cost as much as possible, i.e., given two plays $\pi,\pi^{\prime}$ such that $\mathsf{cost}_{i}(\pi)<\mathsf{cost}_{i}(\pi^{\prime})$, player $i$ prefers $\pi$ to $\pi^{\prime}$. We extend $<$ to tuples of costs as follows: $(\mathsf{cost}_{i}(\pi))_{i\in\mathcal{P}}<(\mathsf{cost}_{i}(\pi^{\prime}))_{i\in\mathcal{P}}$ if $\mathsf{cost}_{i}(\pi)\leq\mathsf{cost}_{i}(\pi^{\prime})$ for all $i\in\mathcal{P}$ and there exists some $j\in\mathcal{P}$ such that $\mathsf{cost}_{j}(\pi)<\mathsf{cost}_{j}(\pi^{\prime})$. Given a play $\pi$, we denote by $\mathsf{Visit}(\pi)$ the set of players $i$ such that $\pi$ visits $T_{i}$, i.e., $\mathsf{Visit}(\pi)=\{i\in\mathcal{P}\mid\mathsf{cost}_{i}(\pi)<+\infty\}$. When for all $i\in\mathcal{P}$ and $e\in E$, $w_{i}(e)=0$, we speak of qualitative reachability games, since $\mathsf{cost}_{i}(\pi)=0$ if $\mathsf{Occ}({\pi})\cap T_{i}\neq\emptyset$ and +$\infty$+ ∞ otherwise.
Strategies and Mealy Machines.
Let $\mathcal{A}{}=(V,E,\mathcal{P},(V_{i})_{i\in\mathcal{P}})$ be an arena. A strategy $\sigma_{i}:\mathsf{Hist}_{i}\rightarrow V$ for player $i$ maps any history $h\in\mathsf{Hist}_{i}$ to a vertex $v\in\mathsf{succ}({\mathsf{last}({h})})$, which is the next vertex that player $i$ chooses to move to after reaching the last vertex in $h$. The set of all strategies of player $i$ is denoted $\Sigma_{i}$. A play $\pi=\pi_{0}\pi_{1}\dots$ is consistent with $\sigma_{i}$ if $\pi_{k+1}=\sigma_{i}(\pi_{0}\dots\pi_{k})$ for all $k\in\mathbb{N}$ such that $\pi_{k}\in V_{i}$. Consistency is naturally extended to histories. A tuple of strategies $\sigma=(\sigma_{i})_{i\in\mathcal{P}}$ with $\sigma_{i}\in\Sigma_{i}$, is called a strategy profile. In an arena initialized at $v_{0}$, we limit the domain of each strategy $\sigma_{i}$ to $\mathsf{Hist}_{i}(v_{0})$; the play $\pi$ starting from $v_{0}$ and consistent with each $\sigma_{i}$ is denoted $\langle\sigma\rangle_{v_{0}}$ and called outcome.
Given an initialized arena $\mathcal{A}$, we can encode a strategy or a set of strategies by a (finite) nondeterministic Mealy machine [DBLP:conf/mfcs/BriceRB23, Pareto-Rational-Verification] $\mathcal{M}=(M,m_{0},\delta,\tau)$ on $\mathcal{A}$, where $M$ is a finite set of memory states, $m_{0}\in M$ is the initial state, $\delta:M\times V\rightarrow 2^{M}$ is the update function, and $\tau:M\times V_{i}\rightarrow 2^{V}$ is the next-move function. Such a machine embeds a (possibly infinite) set of strategies $\sigma_{i}$ for player $i$, called compatible strategies. Formally, $\sigma_{i}$ is compatible with $\mathcal{M}$ if there exists a mapping $h\mapsto m_{h}$ such that $m_{hv}\in\delta(m_{h},v)$ for every $hv\in\mathsf{Hist}(v_{0})$ (with $m_{h}=m_{0}$ when $h$ is empty), and when $v\in V_{i}$, $\sigma_{i}(hv)\in\tau(m_{h},v)$. An example of such a machine $\mathcal{M}$ is given in LABEL:appendix:exMealyMachine. We denote by $\llbracket\mathcal{M}\rrbracket$ the set of all strategies compatible with $\mathcal{M}$. The memory size of $\mathcal{M}$ is equal to $|M|$. We say that $\mathcal{M}$ is deterministic when the image of both functions $\delta$ and $\tau$ is a singleton. Thus when $\mathcal{M}$ is deterministic, =$\llbracket\mathcal{M}\rrbracket=\{\sigma_{i}\}$ and $\sigma_{i}$ is called finite-memory, and when additionally $|M|=1$, $\sigma_{i}$ is called memoryless.
3 Studied Problems
In this section, within the context of rational synthesis and verification, we consider a reachability game $\mathcal{G}=(\mathcal{A},(T_{i})_{i\in\mathcal{P}})$ with $\mathcal{A}$ an initialized weighted arena and $\mathcal{P}=\{0,1,\dots,t\}$ such that player 00 is a specific player, often called system or leader, and the other players $1,\dots,t$ compose the environment and are called followers. Player 00 announces his strategy $\sigma_{0}$ at the beginning of the game and is not allowed to change it according to the behavior of the other players. The response of those players to $\sigma_{0}$ is supposed to be rational, where the rationality can be described as the outcome of a Nash equilibrium [Nash50] or as a Pareto-optimal play [Stackelberg-Pareto-Synthesis].
Nash Equilibria.
A strategy profile for the environment is a Nash equilibrium if no player has an incentive to unilaterally deviate from this profile. In other words, no player can improve his cost by switching to a different strategy, assuming that the other players stick to their current strategies. Formally, given the initial vertex $v_{0}$ and a strategy $\sigma_{0}$ announced by player 00, a strategy profile $\sigma=(\sigma_{i})_{i\in\mathcal{P}}$ is called a 00-fixed Nash equilibrium (00-fixed NE) if for every player $i\in\mathcal{P}\!\setminus\!\{0\}$ and every strategy $\tau_{i}\in\Sigma_{i}$, it holds that $\mathsf{cost}_{i}(\langle\sigma\rangle_{v_{0}})\leq\mathsf{cost}_{i}(\langle\tau_{i},\sigma_{-i}\rangle_{v_{0}})$, where $\sigma_{-i}$ denotes $(\sigma_{j})_{j\in\mathcal{P}\setminus\{i\}}$, i.e., $\tau_{i}$ is not a profitable deviation. We also say that $\sigma$ is a $\sigma_{0}$-fixed NE to emphasize the strategy $\sigma_{0}$ of player 00.
Pareto-Optimality.
When all players collaborate to obtain a best cost for everyone, we need another concept of rationality. In that case, we suppose that the players in $\mathcal{P}\!\setminus\!\{0\}$ form a single player, player $1$, that has a tuple of targets sets $(T_{i})_{i\in\{1,\dots,t\}}$. For each play $\pi\in\mathsf{Plays}(v_{0})$, player $1$ gets a cost tuple $\mathsf{cost}_{\textnormal{env}}({\pi})=(\mathsf{cost}_{i}(\pi))_{i\in\{1,\dots,t\}}$, and prefers $\pi$ to $\pi^{\prime}$ if $\mathsf{cost}_{\textnormal{env}}({\pi})<\mathsf{cost}_{\textnormal{env}}({\pi^{\prime}})$ for the componentwise partial order <<< over ({+}$(\mathbb{N}\cup\{+\infty\})^{t}$. Given such a modified game and a strategy $\sigma_{0}$ announced by player 00, we consider the set $C_{\sigma_{0}}$ of cost tuples of plays consistent with $\sigma_{0}$ that are Pareto-optimal for player $1$, i.e., minimal with respect to $<$. Hence, $C_{\sigma_{0}}=\min\{\mathsf{cost}_{\textnormal{env}}({\pi})\mid\pi\in\mathsf{Plays}(v_{0})\text{ consistent with }\sigma_{0}\}$. Notice that $C_{\sigma_{0}}$ is an antichain. A cost tuple $p$ (called cost in the sequel) is said to be $\sigma_{0}$-fixed Pareto-optimal ($\sigma_{0}$-fixed PO or simply 00-fixed PO) if $p\in C_{\sigma_{0}}$. Similarly, a play is said to be $\sigma_{0}$-fixed PO if its cost is $\sigma_{0}$-fixed PO.
In some problems studied in this paper, we will have to consider games such that all vertices owned by player 00 have only one successor, which means that player 00 has no choice but to choose this successor. In this case, we say that player $1$ is the only one to play.
Rational Verification.
We now present the studied decision problems related to the concept of rational verification. Given some threshold $c\in\mathbb{N}$, the goal is to verify that a strategy $\sigma_{0}$ announced by player 00 guarantees him a cost $\mathsf{cost}_{0}(\pi)\leq c$ whatever the rational response $\pi$ of the environment. By rational response, we mean either a $\sigma_{0}$-fixed NE outcome $\pi$, or a $\sigma_{0}$-fixed PO play $\pi$. The strategy $\sigma_{0}$ is usually given as a deterministic Mealy machine. We can go further: with a nondeterministic Mealy machine, we want to verify whether all strategies $\sigma_{0}\in\llbracket\mathcal{M}\rrbracket$ are solutions. In the latter case, we speak about universal verification.
Problem 1.
Given a reachability game $\mathcal{G}$ with an initialized arena, a nondeterministic Mealy machine $\mathcal{M}_{0}$ for player 00, and a threshold $c\in\mathbb{N}$,
  • If $\llbracket\mathcal{M}_{0}\rrbracket=\{\sigma_{0}\}$, the Non-Cooperative Nash Verification problem ($\mathsf{NCNV}$) asks whether for all $\sigma_{0}$-fixed NEs $\sigma$, it holds that $\mathsf{cost}_{0}(\langle\sigma\rangle_{v_{0}})\leq c$.
  • The Universal Non-Cooperative Nash Verification problem ($\mathsf{UNCNV}$) asks whether for all $\sigma_{0}\in\llbracket\mathcal{M}_{0}\rrbracket$ and all $\sigma_{0}$-fixed NEs $\sigma$, it holds that $\mathsf{cost}_{0}(\langle\sigma\rangle_{v_{0}})\leq c$.
  • If $\llbracket\mathcal{M}_{0}\rrbracket=\{\sigma_{0}\}$, the Non-Cooperative Pareto Verification problem ($\mathsf{NCPV}$) asks whether for all $\sigma_{0}$-fixed PO plays $\pi$, it holds that $\mathsf{cost}_{0}(\pi)\leq c$.
  • The Universal Non-Cooperative Pareto Verification problem ($\mathsf{UNCPV}$) asks whether for all $\sigma_{0}\in\llbracket\mathcal{M}_{0}\rrbracket$ and all $\sigma_{0}$-fixed PO plays $\pi$, it holds that $\mathsf{cost}_{0}(\pi)\leq c$.
Rational Synthesis.
We consider the more challenging problem of rational synthesis. Given a threshold $c\in\mathbb{N}$, the goal is to synthesize a strategy $\sigma_{0}$ for player 00 (instead of verifying some $\sigma_{0}$) that guarantees him a cost $\mathsf{cost}_{0}(\pi)\leq c$ whatever the rational response $\pi$ of the environment. We also consider the simpler problem where the environment cooperates with the leader by proposing some rational response $\pi$ that guarantees him a cost $\mathsf{cost}_{0}(\pi)\leq c$.
Problem 2.
Given a reachability game $\mathcal{G}$ with an initialized arena and a threshold $c\in\mathbb{N}$,
  • The Cooperative Nash Synthesis ($\mathsf{CNS}$) problem asks whether there exists $\sigma_{0}\in\Sigma_{0}$ and a $\sigma_{0}$-fixed NE $\sigma$ such that $\mathsf{cost}_{0}(\langle\sigma\rangle_{v_{0}})\leq c$.
  • The Non-Cooperative Nash Synthesis ($\mathsf{NCNS}$) problem asks whether there exists $\sigma_{0}\in\Sigma_{0}$ such that for all $\sigma_{0}$-fixed NEs $\sigma$, it holds that $\mathsf{cost}_{0}(\langle\sigma\rangle_{v_{0}})\leq c$.
  • The Cooperative Pareto Synthesis ($\mathsf{CPS}$) problem asks whether there exists $\sigma_{0}\in\Sigma_{0}$ and a $\sigma_{0}$-fixed PO play $\pi$ such that $\mathsf{cost}_{0}(\pi)\leq c$.
  • The Non-Cooperative Pareto Synthesis ($\mathsf{NCPS}$) problem asks whether there exists $\sigma_{0}\in\Sigma_{0}$ such that for all $\sigma_{0}$-fixed PO plays $\pi$, it holds that $\mathsf{cost}_{0}(\pi)\leq c$.
[…]
Figure 1: An example illustrating the two concepts of rational response.
$v_{0}$$v_{1}$$(1,0)$$(1,0)$$(1,0)$
Figure 2: An example showing that PO lasso plays in the $\mathsf{coNCPV}$ problem may have an exponential length.
Example 3.1.
To illustrate these problems, let us study a simple example depicted in Figure 2 with three players: the system, player 00, and two players in the environment, players $\square$ and $\diamond$. Player 00 owns the circle vertices, player $\square$ owns the square initial vertex $v_{0}$, and player $\diamond$ owns the diamond vertex $v_{2}$. Each player $i$ has a target set, $T_{0}=\{v_{3},v_{4}\}$, $T_{\text{\scriptsize$\square$\normalsize}}=\{v_{3},v_{5}\}$ and $T_{\text{\large$\diamond$\normalsize}}=\{v_{1},v_{4}\}$, and a constant weight $w_{i}(e)=1$ for all $e\in E$. When a vertex $v$ is in $T_{i}$, we depict the symbol of player $i$ nearby $v$. As the graph is acyclic, the possible player strategies are all memoryless. In the sequel, we thus only indicate the successor chosen by the player.
Let us show that $\sigma_{0}$ defined by $\sigma_{0}(v_{1})=v_{2}$ is a solution to the
Main Results.
Our main results for Problems 1-2 are the following ones when the rational responses of the environment are 00-fixed PO plays. One problem was already solved in [gaspard-pareto-quanti-reach].
Theorem 3.2.
  1. (a)
    The Non-Cooperative Pareto Verification problem is $\Pi_{2}^{\mathsf{P}}$-complete.
  2. (b)
    The Universal Non-Cooperative Pareto Verification problem is $\mathsf{PSPACE}$-complete.
  3. (c)
    The Cooperative Pareto Synthesis problem is $\mathsf{PSPACE}$-complete.
  4. (d)
    The Non-Cooperative Pareto Synthesis problem is $\mathsf{NEXPTIME}$-complete [gaspard-pareto-quanti-reach].
For 00-fixed NE responses of the environment, we obtain the next main results.
Theorem 3.3.
  1. (a)
    The Non-Cooperative Nash Verification problem is $\mathsf{coNP}$-complete.
  2. (b)
    The Universal Non-Cooperative Nash Verification problem is $\mathsf{coNP}$-complete.
  3. (c)
    The Cooperative Nash Synthesis problem is $\mathsf{NP}$-complete.
  4. (d)
    The Non-Cooperative Nash Synthesis problem is $\mathsf{EXPTIME}$-hard, already with a two-player environment. With a one-player environment, it is in $\mathsf{EXPTIME}$ and $\mathsf{PSPACE}$-hard.
These complexity results depend on the size $|V|$ of the arena, the number $t$ of players $i$ (resp. target sets $T_{i}$) in case of 00-fixed NE responses (resp. 00-fixed PO responses), the maximal weight $W$ encoded in binary appearing in the functions $w_{i}$, the threshold $c$ encoded in binary, and the size $|M|$ of the Mealy machine $\mathcal{M}_{0}$ (for the verification problems). Note that for all problems except the NCNS problem, the complexity classes are the same for both qualitative and quantitative frameworks (see Table 1). Hence, in the case of a unary encoding of the weights and the threshold $c$, we get the same complexity classes. Only the most challenging proofs are provided in the paper, while the other proofs or results derived from the literature are deferred in the Appendix.
In this paper, we focus on zero or positive weights, because with negative weights, there are simple examples of one-player games with no NE or no PO plays (thus with no rational responses). Furthermore, considering any weights leads to the undecidability of the
Problem 6.
NCNS and